OpenStudy (anonymous):

What is the possible value for the missing term of the geometric sequence? 37, , 592 a.148 b.4 c.41 d.2368

OpenStudy (anonymous):

Let the missing term be x So, x/37 = 592/x x^2 = 37 x 592 = 21904 x = 148